OpenStudy (anonymous):

helppppppppppppppppp 10 medals

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) [6.04] MA.7.G.4.4 How many centigrams are in one hundred milligrams? 1 centigram ten centigrams one thousand centigrams ten thousand centigrams Question 2 (Multiple Choice Worth 1 points) [6.02] MA.7.A.1.3 Triangle L Q H is similar to triangle M B Z. Which of the following must be true? triangle L Q H must be congruent to triangle M B Z. line segment Q H is congruent to line segment B Z If triangle L Q H is an obtuse triangle, triangle M B Z is an obtuse triangle. line segment L Q is congruent to line segment Z M

OpenStudy (mathstudent55):

\(100~mg \times \dfrac{1~g}{1000~mg} \times \dfrac{100 ~cg}{1~g} \)

OpenStudy (mathstudent55):

Multiply it out.

OpenStudy (mathstudent55):

When two triangles are congruent, the three angles of one triangle are congruent to the three corresponding angles of the other triangle. Corresponding sides may or may not be congruent.